# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ 4.1 Relationship to Hamiltonian learning
with $U=e^{itH(\bm{\lambda})}$, where the evolution time $t$ is known to us. In [HKT22a] the authors provide a classical algorithm which learns the unknown $\bm{\lambda}$ from the expectation values of the Pauli string $P_{i}$ on polynomially many copies of the evolved state $U\rho U^{\dagger}$, for particular choices of initial states $\rho$’s. We can rephrase the Hamiltonian learning task in terms of a learning problem with many similarities to the ones we considered before. Concretely we define the corresponding concept class
$\mathcal{F}_{\bm{\lambda}}$ =$=\{f^{\bm{\lambda}}(\bm{x})\in\mathbb{R}\;\;|\;\;\bm{\lambda}\in[-1,1]\}$ (12)
with $\text{with }\;f^{\bm{\lambda}}(\bm{x}):$ $\;x\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho(\bm{x})O(\bm{\lambda})]\;\;\&\;\;O(\bm{\lambda})=\sum_{i}^{m}U(\bm{\lambda})^{\dagger}P_{i}U(\bm{\lambda}).$ (13)
@@
where $x$ describes the input state $\rho(\bm{x})$, and $U=e^{itH(\bm{\lambda})}$. Considering $\mathcal{T}^{\bm{\lambda}}={(\bm{x}_{\ell},\text{Tr}[\rho(\bm{x}_{\ell})O(\bm{\lambda})])}_{\ell}$ as training data, the connection to the Hamiltonian learning task of [HKT22a] becomes evident if we demand that the learning algorithm needs only to identify, rather than evaluate, the correct concept that generated the data. It is worth noting that the concept class $\mathcal{F}_{\bm{\lambda}}$ closely resembles the concept class for the time evolution problem described earlier, for which a learning advantage was demonstrated. The main difference is that in $\mathcal{F}_{\bm{\lambda}}$ the unknown observable is no longer a linear combination of Pauli strings, instead each Pauli string in $O(\bm{\lambda})$ is parameterized by a unitary. As we are requiring to only identify the correct concept and considering that the authors in [HKT22a] provide an algorithm capable of solving the Hamiltonian Learning problem in polynomial time (for sufficiently small $t$), one might suppose that the classical difficulty associated with the learning advantage presented in this work arises solely from the challenge of evaluating the correct concept rather than from the task of identifying it. This would answer another question left open in [GD23], at least in the case of concepts labeled by observables with an unknown unitary parameterization. However it must be noted that for the Hamiltonian learning algorithm in [HKT22a] to work, the input states $\rho(\bm{x})$ in Eq.(12) are of a very particular and simple form and do not come from a distribution $\mathcal{D}$ of $\mathrm{BQP}$ complete quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ as we assumed in our classical hardness result. For general input distributions, the hardness of identification thus remains elusive. In this regard it is important to note that the task of identification can take various forms. For instance, here we are focusing on identifying the concept within the concept class that generated the data, and in this case, it is generally unclear when this setting allows a classical-efficient solution. In a more general scenario one could ask if the learning algorithm is able to identify a model from a hypothesis class that differs from the concept class but that it is still capable of achieving the learning condition. Colloquially here the question becomes: can traditional machine learning determine which quantum circuit could accurately label the data, even though classical computers lack the capability to evaluate such quantum circuits. In [GD23] it was proven that in this broader context provable speed-ups in identification are no longer possible. This is because the classical algorithm can successfully identify a sophisticated quantum circuit that carries out the entire quantum machine learning process as its subroutine, thereby delegating the learning aspect to the labeling function itself, prior to evaluating a data point. Thus the task of identification is indeed only interesting when the hypothesis class is somehow restricted, which is often the case in practically relevant learning scenarios.
@@ Acknowledgments
Appendix A Proof of Theorem 1
As anticipated in the main text, the proof of the main theorem of our work consists of two parts. First, we show that there exists a family of hamiltonians $H_{\mathrm{hard}}\in\mathcal{H}_{\mathrm{hard}}$ and observables $O$ such that the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ is not classically learnable, unless $\mathrm{BQP}\subseteq\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ for every distribution $\mathcal{D}$. Second, we rigorously prove that Algorithm 1 learns $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ in polynomial time. We formally derive the proofs for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ of measurements on time-evolved states. However, as we discussed in the main text, the same proofs can be easily extended for the class $\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}$ of measurements on ground state of local Hamiltonians through the Kitaev’s circuit-to-hamiltonian construction. Before stating the main proof, let prove an useful lemma which guarantees the BQP-completeness of constant time Hamiltonian evolution following the idea in present in [Chi04].
Lemma 3 (Constant time evolution is BQP-complete).
@@ Appendix A Proof of Theorem 1
Consider the Feymann clock Hamiltonian [Fey85, Nag10]:
@@
We now follow the idea in [Chi04] and modify the Hamiltonian in (15) in the following way:
@@
The idea in [Chi04] is to associate each state in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ to a quantum system of total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ with $z$ component $j-\frac{k}{2}$. The association is possible by the fact that a system with total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ will allow $k+1$ states with $z$ components of values $-\frac{k}{2},-\frac{k}{2}+1,...,\frac{k}{2}-1,\frac{k}{2}$. These are exactly the $k+1$ states in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ and it is possible to move among them defining the corresponding ladder operators :
@@
A.1 Classical hardness
Theorem 4 (Lemma 1 in the main text).
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated input distribution $\mathcal{D}$ and a Hamiltonian $H_{\mathrm{hard}}$ such that no randomized polynomial-time classical algorithm $A_{c}$ is such that $h_{cl}(.)=A_{c}(T^{\bm{\alpha}},.)$ satisfies the learning condition of Def. 1 for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with $x$ sampled from $\mathcal{D}$.

### v2

@@ V.2 Relationship to Hamiltonian learning
with $U=e^{itH(\bm{\lambda})}$, where the evolution time $t$ is known to us. In [35] the authors provide a classical algorithm which learns the unknown $\bm{\lambda}$ from the expectation values of the Pauli string $P_{i}$ on polynomially many copies of the evolved state $U\rho U^{\dagger}$, for particular choices of initial states $\rho$’s. We can rephrase the Hamiltonian learning task in terms of a learning problem with many similarities to the ones we considered before. Concretely we define the corresponding concept class
$\mathcal{F}_{\bm{\lambda}}=\{f^{\bm{\lambda}}(\bm{x})\in\mathbb{R}\;\;|\;\;\bm{\lambda}\in[-1,1]^{m}\}$ (11)
with: $\;\;f^{\bm{\lambda}}(\bm{x}):\;\bm{x}\in\mathcal{X}\subseteq\{0,1\}^{n}\rightarrow\text{Tr}[\rho(\bm{x})O(\bm{\lambda})]$
$O(\bm{\lambda})=\sum_{i=1}^{m}U(\bm{\lambda})^{\dagger}P_{i}U(\bm{\lambda}).$
@@
where $\bm{x}$ describes the input state $\rho(\bm{x})$, and $U=e^{itH(\bm{\lambda})}$. We note that the specific choice of the observable $O(\bm{\lambda})=\sum_{i=1}^{m}U(\bm{\lambda})^{\dagger}P_{i}U(\bm{\lambda})$ is made to align with the work in [35], where the learning algorithm solves the task by directly measuring the Pauli operators $\{P_{i}\}_{i}$ after evolving the input states with $U=e^{itH(\bm{\lambda})}$. In this sense, by considering $\mathcal{T}^{\bm{\lambda}}={(\bm{x}_{\ell},\text{Tr}[\rho(\bm{x}_{\ell})O(\bm{\lambda})])}_{\ell}$ as training data, the connection to the Hamiltonian learning task of [35] becomes evident if we demand that the learning algorithm needs only to identify, rather than evaluate, the correct concept that generated the data. It is worth noting that the concept class $\mathcal{F}_{\bm{\lambda}}$ closely resembles the concept class for the time-evolution problem described earlier, for which a learning advantage was demonstrated. The main difference is that in $\mathcal{F}_{\bm{\lambda}}$ the unknown observable is no longer a linear combination of Pauli strings, instead each Pauli string in $O(\bm{\lambda})$ is parameterized by a unitary. As we are requiring to only identify the correct concept and considering that the authors in [35] provide an algorithm capable of solving the Hamiltonian Learning problem in polynomial time (for sufficiently small $t$), one might suppose that the classical difficulty associated with the learning advantage presented in this work arises solely from the challenge of evaluating the correct concept rather than from the task of identifying it. This would answer another question left open in [11], at least in the case of concepts labeled by observables with an unknown unitary parameterization. However it must be noted that for the Hamiltonian learning algorithm in [35] to work, the input states $\rho(\bm{x})$ in Eq.(11) are of a very particular and simple form and do not come from a distribution $\mathcal{D}$ of $\operatorname{\mathsf{BQP}}$-complete quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ as we assumed in our classical hardness result. For general input distributions, the hardness of identification thus remains elusive. In this regard it is important to note that the task of identification can take various forms. For instance, here we are focusing on identifying the concept within the concept class that generated the data, and in this case, it is generally unclear when this setting allows a classical-efficient solution. In a more general scenario one could ask if the learning algorithm is able to identify a model from a hypothesis class that differs from the concept class but that it is still capable of achieving the learning condition. Colloquially here the question becomes: can traditional machine learning determine which quantum circuit could accurately label the data, even though classical computers lack the capability to evaluate such quantum circuits. In [11] it was proven that in this broader context provable speed-ups in identification are no longer possible. This is because the classical algorithm can successfully identify a sophisticated quantum circuit that carries out the entire quantum machine learning process as its subroutine, thereby delegating the learning aspect to the labeling function itself, prior to evaluating a data point. Thus the task of identification is indeed only interesting when the hypothesis class is somehow restricted, which is often the case in practically relevant learning scenarios.
VI Discussion
In this paper, we propose a machine learning task that demonstrates a quantum learning advantage under the assumption that quantum processes cannot be simulated by polynomial-size classical circuits (i.e., $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$). The learning problem involves reproducing a partially unknown observable from classical, measured-out data. While classical learning theory guarantees the computational hardness of the task for input states originating from certain quantum processes, we have developed a quantum learning algorithm capable of solving it for any input distribution. In this discussion, we consider whether this abstract learning problem can reasonably model physically relevant scenarios. We identify three key components for our framework: i) the preparation of a “hard” quantum state from a large family, indeed ensuring that $U$ specifies a classically-hard but quantum realizable operation  [^5] For our purposes, it is sufficient (though not necessary) to have a parameterized heralded state $\rho(x)$ such that measuring an observable $O$ on it produces the function $x\mapsto\mathrm{Tr}[\rho(x)O]$, which cannot be implemented using polynomial-sized circuits. (see below for details); ii) the heralded yet uncontrolled nature of the selection of $U(x)$, reflecting the fact that data is generated under an unknown distribution over $x$ in the PAC learning framework; and iii) the task of learning unknown observables itself.
We next discuss the extent to which these three criteria correspond to natural learning tasks.
  1. i)
    Hard state preparation. While in our proof we deal with the contrived Kitaev Hamiltonian, this is only for technical simplicity. In general, all that is required is a scenario where a family of parametrized “hard” states (|$\ket{\psi(x)}=U(x)\ket{0}$) is measured, and where there is strong evidence to suggest that computing the map $x\mapsto\bra{\psi(x)}O\ket{\psi(x)}$ (where $x$ is the description of the state in the family) is beyond the capabilities of classical polynomial-size circuits (i.e., not in $\mathsf{P/poly}$ for the decision variant of the problem). Time-evolution scenarios (where $U(x)$ represents the time-evolution of a time-dependent or independent Hamiltonian) and ground-state scenarios (where $U(x)\ket{0}$ is the desired ground state) are natural candidates for such state preparation. To provide the strongest arguments that the resulting map is also classically hard, it is useful to work with Hamiltonians known to be computationally hard (i.e. $\operatorname{\mathsf{BQP}}$- or $\mathsf{QMA}$- hard in the context of time-evolution or ground-state problems, respectively). With these conditions in mind, we can still identify numerous previously recognized families of physical Hamiltonians that satisfy these criteria. For time-evolution, to list a few known cases of $\operatorname{\mathsf{BQP}}$-complete Hamiltonians and otherwise specified evolutions which give raise to $\operatorname{\mathsf{BQP}}$-complete problems:
    • $\bullet$
      Various variants of the Bose-Hubbard model [37].
    • $\bullet$
      Stoquastic Hamiltonians [38].
    • $\bullet$
      Ferromagnetic Heisenberg model [37].
    • $\bullet$
      The $XY$ model [39].
    • $\bullet$
      Estimating the scattering probabilities in massive field theories [40], (more precisely, estimating the vacuum-to-vacuum transition amplitude, in the presence of spacetime-dependent classical sources, for a massive scalar field theory in $(1+1)$ dimensions).
    • $\bullet$
      Simulation of topological quantum field theories [41].
    For the ground state version, $\mathsf{QMA}$ hardness results are known for the following Hamiltonians (the first two examples are actually in $\mathsf{MA}$, a subset of $\mathsf{QMA}$):
    • $\bullet$
      Stoquastic local Hamiltonians [42].
    • $\bullet$
      Bose-Hubbard with negative hoppings [42].
    • $\bullet$
      Bose-Hubbard with poistive hoppings [43].
    • $\bullet$
      Antiferromagnetic Heisenberg model [39].
    • $\bullet$
      The $XY$ model [39].
    • $\bullet$
      Electronic structure problem [44] for quantum chemistry.
    • $\bullet$
      Certain supersymmetric quantum mechanical theories [45].
    We note that the list above is, to our knowledge, not exhaustive. Focusing on the ground state setting, the computational hardness involved in preparing these states is reflected in our quantum learning algorithm, however, we do not imply that either nature (in the realization of the dataset) or our quantum algorithm is solving $\mathsf{QMA}$-hard problems. In practical scenarios, any $\mathsf{QMA}$ problem can include instances that are solvable within $\mathsf{BQP}$ (as indeed $\operatorname{\mathsf{BQP}}$ is contained in $\mathsf{QMA}$, and so all $\operatorname{\mathsf{BQP}}$ problems correspond to some instances of $\mathsf{QMA}$-hard problems). For those distributions over the inputs, which are arguably the ones realized in nature, there can be efficient quantum algorithms. Specifically, while the Hamiltonian families listed above define a broader class of physically relevant systems, the instances encountered in real-world scenarios typically fall within easier subclasses that are likely to involve $\mathsf{BQP}$-hard problems. This aligns well with our framework. An example of a setting where the quantum learning algorithm is easy to specify occurs in systems where classical approximations of ground states (such as mean-field solutions, matrix product approximations, etc.) are not exponentially far from the true ground states. These systems, while belonging to $\mathsf{QMA}$-like families, could result in $\operatorname{\mathsf{BQP}}$-complete problems (by virtue of the guided Hamiltonian problem construction [46]). Other examples of quantumly efficient simulable processes involve Hamiltonians connected by an inverse-polynomially gapped adiabatic path, which can occur in critical Hamiltonians or near phase transition boundaries, with the appropriate scaling of system size. In all of these cases, our approach not only provides a natural learning scenario but also an efficient quantum learning algorithm that effectively leverages a guiding state. Finally, as we transition to practical applications, it is important to note that the emphasis on $\operatorname{\mathsf{BQP}}$-complete problems is primarily relevant when aiming for formal proofs of asymptotic advantages. In particular, our separation results holds for any quantum process which maps $x\rightarrow\rho(x)$ and for which there exists an observable $O$ such that the function: $x\rightarrow\operatorname{tr}[\rho(x)O]$ is outside $\mathsf{P/poly}$, thus contained in the class $(\mathsf{P/poly})^{\mathrm{c}}\cap\operatorname{\mathsf{BQP}}$ 666$(\mathsf{P/poly})^{\mathrm{c}}$ indicates the complement of the class $\mathsf{P/poly}$. ($\subseteq\operatorname{\mathsf{BQP}}$-complete). Many systems of interest may not be $\operatorname{\mathsf{BQP}}$-complete but are likely beyond the reach of classical poly-sized circuits. Potential examples may include problems in glassy systems, Hamiltonians related to exotic highly correlated states of matter (e.g., spin liquids), highly correlated molecules (such as those involving heavier metals), and certain areas of high-energy physics, like quantum chromodynamics. While many of these problems may not be strictly $\operatorname{\mathsf{BQP}}$-complete, they may remain intractable for classical computers and could potentially be addressed by quantum ones, making them strong candidates for demonstrating quantum advantages in our learning task. We acknowledge, however, that the hardness of these last examples listed above remains purely conjectural and has yet to be more precisely assessed.
  2. ii)
    Data distribution: heralding without full control. In the formulation of our learning problems, we assume that the input Hamiltonians, or more generally the resulting states, are drawn from a distribution. While it may not be unusual to have no information about which Hamiltonian is acting on the system, our framework assumes that this information is also revealed in the data point $x$. Specifically, for our approach to work, we require both that the Hamiltonian is chosen randomly and that we are later informed about which Hamiltonian was selected—essentially heralding without control  [^7] If the Hamiltonian were fully controlled, we would no longer be in a PAC learning scenario, but something closer to Angluin’s query model [55]. While this setup would also be interesting, the learning objectives would differ.. Such scenarios can occur in situations where control is limited, but some type of partial (process) tomography allows the reconstruction of the implemented Hamiltonian afterward. For example, in fields like materials science and condensed matter physics, we frequently do not have full control over the conditions affecting a sample, but we can analyze it afterward. This is particularly true in cases like material doping or the introduction of nitrogen-vacancy (NV) centers. In condensed matter systems, doping introduces additional electrons or holes into a material. However, the precise concentration and spatial distribution of dopants are challenging to control at the atomic scale, leading to variability in the resulting electronic structure [49]. Similarly, lattice imperfections, such as strain, dislocations, or defects, perturb key Hamiltonian parameters like hopping amplitudes and interaction strengths, often in unpredictable ways [50]. Impurities, including unintentional substitutions or vacancies, modify on-site energies and introduce disorder [51]. Collectively, these factors result in effective Hamiltonians that deviate from the intended design, necessitating a statistical approach to account for variability. Such challenges are not unique to condensed matter systems. In quantum optics, heralded operations used to generate entangled photon pairs rely on probabilistic detection events, meaning the effective realized quantum operation depends on external factors such as photon loss or background noise. Similarly, in high-energy physics, the governing Hamiltonian of particle collisions is subject to environmental factors like background radiation or fluctuating particle flux, making precise control impractical. In both cases, experimentalists work within a probabilistic framework, relying on statistical inference rather than attempting direct control of the Hamiltonian. Our model reflects these realities by treating input Hamiltonians as drawn from an underlying distribution. This approach is well-suited to systems with intrinsic or practical variability, where repeated experiments often produce diverse realizations of effective Hamiltonians. By adopting a distributional perspective, we capture the statistical nature of Hamiltonian variability and address learning problems under realistic experimental constraints. Furthermore, since our work requires learnability for any input distribution, rather than focusing on specific distributions as in [11], we more accurately address the case where the underlying distributions are unknown. This makes our framework further aligned with real-world settings than previous approaches.
  3. iii)
    Learning unknown observables. There are many natural scenarios where the actual measurement performed is not fully characterized or is entirely unknown. One clear example is learning order parameters (as observables) based on, e.g., data labeled using indirect indicators, such as entanglement entropy, which are not simple linear functions of the state. A similar scenario occurs when the order parameter is known, but the corresponding observable cannot be directly implemented on the given device. In these cases, a learning model can infer a proxy for the true order parameter as a linear combination of observables that are feasible to implement. Additionally, since time-evolution under a parametrized Hamiltonian can be considered part of the measurement in the case of unitarily parametrized observables, the scenario of partially unknown observables also includes situations where certain aspects of the dynamics are unknown, while the measurement itself remains fixed. Another relevant setting concern experimental contexts where, due to incomplete device characterization, the measurements realized may not align with the intended ones, reflecting a partially unknown scenario. It is important to note that limiting the degree of ignorance in the setting corresponds to restricting or biasing the concept class structure. In this case, learnability still holds (regardless of the prior) along with classical hardness, as long as there remains at least one “hard” measurement within the concept class. The characterization and learning of observables have been extensively studied, precisely because it represents a plausible scenario, especially within the framework of quantum measurement tomography, which was first introduced in [52]. Specifically, in [52] a setting where unknown observables emerge was identified, when a system interacts with external degrees of freedom before being measured by a fixed observable. This is particularly relevant in cases involving reservoirs or other mechanisms that induce losses and decoherence effects. Subsequent works, such as [53], used the task of learning unknown measurements to characterize experimental detectors. This task is especially pertinent in quantum computing, where measurement devices are often imperfect due to noise, making it crucial to understand and mitigate the effects of these imperfections on quantum measurements.
Combining meaningful settings from these three classes of assumptions results in more realistic learning tasks that can be captured by our abstract learning framework.
@@ Acknowledgments
Appendix A Proof of classical hardness
As anticipated in the main text, in order to prove our main Theorem 1 we need to show both classical hardness and quantum learnability for the time-evolution learning problem. In this section, we prove Theorem 2 for the time-evolution learning problem, formally described through the concept class $\mathcal{F}^{H,O}_{\mathrm{evolved}}$. It is important to notice that, as we discussed in the main text, the same proofs can be easily extended for the class $\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}$ of measurements on ground state of local Hamiltonians through the Kitaev’s circuit-to-hamiltonian construction.
A.1 $\operatorname{\mathsf{BQP}}$-completeness of constant time-evolution
Before stating the proof for Therem 2, let us prove an useful lemma which guarantees the $\operatorname{\mathsf{BQP}}$-completeness of constant time Hamiltonian evolution following the idea in present in [26]. It is important to notice that the Hamiltonians used in Lemma 1 will not have a constant norm. This is however not a problem for our quantum algorithm as their norms will scale only polynomially with the number of qubits.
Lemma 1 (Constant time-evolution is $\operatorname{\mathsf{BQP}}$-complete).
@@ A.1 $\operatorname{\mathsf{BQP}}$-completeness of constant time-evolution
Consider the Feyman clock Hamiltonian [24, 25]:
@@
We now follow the idea in [26] and modify the Hamiltonian in (13) in the following way:
@@
The idea in [26] is to associate each state in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ to a quantum system of total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ with $z$ component $j-\frac{k}{2}$. The association is possible by the fact that a system with total angular momentum $\frac{k}{2}(\frac{k}{2}-1)$ will allow $k+1$ states with $z$ components of values $-\frac{k}{2},-\frac{k}{2}+1,...,\frac{k}{2}-1,\frac{k}{2}$. These are exactly the $k+1$ states in {|$\{\ket{\psi_{j}}\}_{j=0}^{k}$ and it is possible to move among them defining the corresponding ladder operators :
@@
A.2 Proof of Theorem 2
The second ingredient in demonstrating the classical hardness of our learning task, in addition to the $\operatorname{\mathsf{BQP}}$-completeness of time-evolution, is a result from classical learning theory contained in [23]. Here, we will state this result in Lemma 2 and, for clarity and didactical purpose, provide an intuitive explanation of its proof.
Lemma 2 (Learnability implies evaluation - Theorem 7 in  [23]).
Suppose $\mathcal{F}$ is learnable in the sense of Def. 1 with input space $\mathcal{X}_{n}=\{0,1\}^{n}$. Then there exists a polynomial $p$ such that for all the concepts $f\in\mathcal{F}$ of size $s$, there exists a classical circuit of size $p(n,s)$ exactly computing $f$
Proof sketch.
In Lemma 2, the size $s$ of a concept refers to the length of its representation under some encoding scheme. For example, in the case of our concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$, $s\sim\mathrm{poly}(n)$ as our concepts can be implemented by polynomial size quantum circuits. The key result underlying Lemma 2 is another finding from [23], which states the following. If the concept class $\mathcal{F}$ is learnable in the sense of Def. 1, then there exists a learning algorithm that, for any target concept $f\in\mathcal{F}$, produces a hypothesis implementable by a circuit of size $\mathrm{poly}(n,s,\log(1/\epsilon))$, which is $\epsilon$-close to the target concept in the sense of Eq.1. This remarkable result is achieved through a smart application of boosting theory. The proof of Lemma 2 then follows by applying the boosted learning algorithm to a training set containing all input points $x\in\{0,1\}^{n}$ and requiring an exponential precision $\epsilon\leq\frac{1}{2^{n}}$. ∎
We are now ready to state the proof of Theorem 2 on the existence of a Hamiltonian $H_{\mathrm{hard}}$ such that the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ is not classically learnable in the sense of Def. 4.
See 2